togo flew his ultralight plane to a landing field 30 mi away. With the wind, the flight took 3/5h. Returning against the wind, the flight took 5/6h. Find the rate of the plane in still air and the rate of the wind.
Answer by checkley77(12844)

D=RT
30/.6=r
r=50 mph with the wind.
30/(5/6)=r
5r/6=30
5r=30*6
5r=180
r=180/5
r=36 mph against the wind.
(50-36)/2=14/2=7 is the wind speed.
50-7=43 speed of the plane.
36+7=43 ditto.